Nathan plans to travel to Europe. There are a number of options available, but he is considering the
following two plans.
Travel Plan A is a plan that consists of a flight cost of $1200 and then charges a fee of $40 per day
Travel Plan B is a plan that charges $75 per day, with unlimited travel time that has the flight charges built
into the cost. If Nathan wants to travel for 30 days, which of the two plans would be the best option? Support your
answer with evidence.

Answer by mananth(16947)

Travel Plan A is a plan that consists of a flight cost of $1200 and then charges a fee of $40 per day
C= 1200 +40*30
C=2400
Travel Plan B is a plan that charges $75 per day, with unlimited travel time that has the flight charges built into the cost.
If Nathan wants to travel for 30 days,
C=75*30 = 2250
Plan B < Plan A